OpenStudy (anonymous):

5/7÷4/5 A 9/11.B 4/7 C. 25/28 D. 5/14

OpenStudy (anonymous):

\[\frac{ 5 }{ 7 } \div \frac{ 4 }{ 5 } = \frac{ 20 }{ 35 } = \frac{ 4 }{ 7 }\] That's how you'd solve it. All you do is flip the second fraction, then multiply. Also simplify if needed :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

That was wrong